Calculate the pH of a rain cloud that has equilibrated with 1.0 ppm by volume SO2
Answer #1

PSO2 = (1 atm)(1 x 10-6) = 1 x 10-6 atm

[H2SO3] = KH·PSO2 = (1.0 M/atm)(1 x 10-6 atm) =1 x 10-6 M

Reaction: H2SO3 ↔ H+ + HSO3- Ka1 = 0.017

Ka1 = 0.017 = [H++ ][HSO3- ] / [H2SO3] = [H+ ]2 / 1 x 10-6

[H+ ] = [(0.017)(1 x 10-6)]0.5 = 4.123 x 10-4 M

pH = 3.38